OpenStudy (dallaslifebaby01):

A sequence has a second term of 1 and a common difference of -3. The fifth term of the sequence is _____.

rishavraj (rishavraj):

\[a_2 - a_1 = d\] so \[1 - a_1 = -3\] get a_1 then use \[a_n = a_1 + (n - 1)d\] to find the fifth term

OpenStudy (dallaslifebaby01):

so would the 5th term be -5?

rishavraj (rishavraj):

nah nah..... u need to plug n = 5 in the equation so it would be \[a_5 = a_1 + (5 - 1)d\] plug a_1 = 4 and d = -3

OpenStudy (dallaslifebaby01):

ooh i see